Using the Intermediate Value Theorem and a calculator, find an interval of length 0.01 that contains a root of x5−x2+2x+3=0, rounding off interval endpoints to the nearest hundredth.

<x<
1
Expert's answer
2020-12-01T01:28:59-0500
"Solution"

For "f(x) = x^5 - x^2 + 2x + 3";


"f(0) = 3\\

f(-1) = -1 - 1 - 2 + 3 = -1"


Thus, as this is a continuous function, from the Intermediate Value Theorem, we know that, as "f(-1) < 0" and "f(0) > 0", there is a root between "-1" and "0."


While we could use the bisection method, the easiest thing is to find "f(x)" for "x" between "-1" and in increments of ".01"


"f(-0.88) = -0.0621319168000003\\

f(-0.87) = 0.00467907929999978"


Then, the Intermediate Value Theorem tells us, at "f(-0.88) < 0" and "f(-0.87) = 0.00467907929999978", there is a root in the interval "(-0.88, -0.87)"